OpenStudy (anonymous):

Graph the equation using the methode of your choice: 3x - 6y= -6

OpenStudy (anonymous):

First solve for y: 3x-6y=-6 -6y=-6-3x y=1+(1/2)x Therefore, the slope of the line is 1/2 and the y-intercept is 1. 

OpenStudy (anonymous):

Thank you.